OpenStudy (anonymous):

The value of x in the proportion 1/2:2/3 = 3/4:x is 1 4/9 1 7/9 1/4

OpenStudy (yttrium):

\[\frac{ 1 }{ 2 } : \frac{ 2 }{ 3 } = \frac{ 3 }{ 4 }: x\] Hence \[\frac{ 1 }{ 2 }x = \frac{ 6 }{ 12 }\] hence \[x = 1\]

OpenStudy (anonymous):

ask god

OpenStudy (anonymous):